find the equation of the parabola whose vertex is V (-3,3) AND WITH DIRECTRIX x3





Answer by lwsshak3(11628)

find the equation of the parabola whose vertex is V (-3,3) AND WITH DIRECTRIX x=3
***
Given parabola opens leftward:
Its basic form of equation:(y-k)^2=-4p(x-h),(h,k)=coordinates of the vertex
For given parabola:
vertex:(-3,3
axis of symmetry: y=3)
p=6 (distance from vertex to directrix on the axis of symmetry)
4p=24
Equation:
(y-3)^2=-24(x+3)